OpenStudy (anonymous):

Find the unknown dimensions. A plastic storage container has a width of one foot longer than the height. The length is four feet longer than the height. The volume is 36 cubic feet. Find the dimensions of the container. x= height x+4= length x+1 = width Please help!!

OpenStudy (danjs):

Volume = Length x Width x Height

OpenStudy (anonymous):

@DanJS So its Volume = x+4*x+1 *x

OpenStudy (danjs):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (danjs):

set that = to 36, and solve for x, the height, from that you get the other width and length

OpenStudy (anonymous):

ok, so 36= x+4*x+1*x

OpenStudy (danjs):

yes, but those are quantities 36 = (x+4)*(x+1)*x

OpenStudy (danjs):

The way you wrote it , it is the same as - x + 4x + x = 6x, not true

OpenStudy (danjs):

you can solve it conventional ways, but if it is a whole number x as the height, the factors of 36 are: 2 * 3 * 6 = x * (x+1)*(x+4)

OpenStudy (danjs):

that should make it easier

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i sort of get it now

OpenStudy (danjs):

you could multiply all that out and get 36 = x^3 + 5x^2 + 4x and solve that way

OpenStudy (danjs):

or 36 = 2 * 3 * 6 = x*(x+1)*(x+4)

OpenStudy (danjs):

2 * 3 * 6 = x *( x+1)*(x+4) = 2*(2+1)*(2+4)

OpenStudy (danjs):

just because it is a "nice" problem from a book, the answer is a nice whole number

OpenStudy (anonymous):

thanks